OpenStudy (anonymous):

If y varies inversely with x and y = -3 when x = 18, find x when y = -2 -33 27 -12 -6

OpenStudy (studygurl14):

Inverse Variation: \(\large y=\Large\frac{k}{x}\), where k= the constant. y = -3 x = 18 First plug in those values to find the constant. Then plug in y = -2, and the value of k for k, and solve for x.

OpenStudy (anonymous):

so -3= \[-3= \frac{ k }{ 18 }\]

OpenStudy (anonymous):

-54? so my new equation would be \[-2=\frac{ -54 }{ x }\]

OpenStudy (anonymous):

@studygurl14

OpenStudy (anonymous):

27.